find theoritical values of Vb and Ve
VCC 20.0V (RB 470k VBT Q2 2N4124 Vioch:- Videy - VE RE >2.2kg
VCC 20.0V RB 47050 VB Q1: 2N4124 V(de): 10.7 V VE V(de): 10.0 V RE: $2.2ko
Answer #1

consider the Circuit shown Vec 201 VCC-IB RB - VBE IERE=0 Consider here Ie= I By Ic this loop 470k & Ic=BIB DE=(1+8) IB + VBEIB = 20-0.4 470K+ (226)(2:26) 19.3 967.2K IB = 0.0199 ma from this drop across the Base Ve will be Vec-VB 20-VB Те RB 470K 20